OpenStudy (anonymous):

Given the circle having a diameter with endpoints (-2, 1) and (6, 7). The equation would be (x-2)2 + (y-4)2 =

OpenStudy (cwrw238):

length of diameter = sqrt((7-1)^2 + (6--2)^2) = sort 100 = 10 and radius = 5 RHS of equation is r^2 = 5^2 = 25 so answer is (x-2)2 + (y-4)2 = 25